OpenStudy (barreraa):

Evaluate x – 2y if x = –3 and y = –6. (Points : 4) 15 9 –9 –15

OpenStudy (barreraa):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

-3+12=12-3=?

OpenStudy (barreraa):

9 right?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (barreraa):

thanks

Nnesha (nnesha):

this question is same like last one please PLEASE at least give him/her chance to doafirst if he/she got that wrong then explain it again

OpenStudy (alexandervonhumboldt2):

gain she ssays i gave a direct answer. WHERE?